Circular No. 8136

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 K2 Eric Christensen, Lunar and Planetary Laboratory, reports the discov-

ery of a comet (discovery observation below) by the Catalina Sky Survey on CCD images taken with the 0.7-m Schmidt telescope. Following posting on the NEO Confirmation Page, many observers noted the obvious cometary nature of the object on CCD images taken during May 27.1–27.2 UT, including R. Elliot (Fall Creek, WI; coma diameter ∼ 10′′), P. R. Holvorcem and M. Schwartz (near Nogales, AZ; coma diameter ∼ 35′′, with a 30′′ tail in p.a. 106^{o}), J. Young (Table Mountain, CA; 10′′ coma and a very faint 40′′ tail in p.a. 115^{o} with a slight curve halfway along its length to p.a. 130^{o}), and J. McGaha (Tucson, AZ; coma diameter 12′′, with slight nuclear condensation and a 6′′ tail).

2003 UT  α_{2000}  δ_{2000}  m_{1}
May 26.18118  8^{h}20^{m}57.^{s}31  +28^{o}34′00′′.4  14.6

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-K43.

T = 2003 Apr. 11.916 TT  ω = 7.^{o}765
Ω = 88.620 }2000.0
q = 0.70200 AU  i = 10.214

SUPERNOVAE 2003ej, 2003ek, AND 2003el Further to IAUC 8134, T. Matheson, P. Challis, and R. Kirshner report

that a spectrum (range 370–750 nm) of SN 2003ej, obtained by M. Calkins on May 22.16 UT, shows it to be a type-II supernova with a blue continuum and incipient P-Cyg lines of hydrogen. Adopting the NED host-galaxy recession velocity (5090 km/s), the expansion velocity derived from the minimum of the Hβ line is 8800 km/s. The Hα emission line is unusual, showing a flat top with some apparent structure, reminiscent of the early spectral stages of SN 1993J (e.g., Matheson et al. 2000, A.J. 120, 1487). A spectrum of SN 2003ek, obtained by Calkins on May 22.43, shows it to be a type-Ia supernova before maximum. Adopting the NED host-galaxy recession velocity (10804 km/s), the supernova expansion velocity is 12000 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 7 ± 2 days before maximum light. A spectrum of SN 2003el (cf. IAUC 8135), obtained by P. Berlind on May 25.27 UT, shows it to be a type-Ic supernova near maximum.

2003 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green